# Dialogue Response Prefetching Based on Semantic Similarity and Prediction Confidence of Language Model

### v1

@@ 1 Introduction
With the progress of research into speech recognition and dialogue response generation, cascade-type spoken dialogue systems (SDSs) are becoming more widely used. Cascade-based SDSs generate spoken responses following two steps [1]. An automatic speech recognition (ASR) model transcribes a user's speech. A dialogue model generates a response from the transcription. In this sequence, the SDSs generate the response after end-of-sentence/speech (EOS), which is determined by the voice activity detection (VAD) model of ASR. However, this framework causes response delays because the user must wait for the system to generate a response after the user's utterance has ended, which decreases the naturalness of an SDS [2].
Speech response prefetching, which prepares dialogue responses before a user has finished speaking, has been proposed to solve this problem [3, 4]. If prefetching is successful, the system can respond immediately after detecting EOS. One of the current response prefetching procedures assumes incremental speech recognition and uses a language model to predict complete user utterances from partially recognized utterances. The language model used to predict complete user utterances is called the prediction model. It addresses the delay problem by preparing the system's speech response in advance for the predicted complete user utterance. The system uses a language model to predict the confidence score for the predicted complete utterance to determine the prefetching timing. The language model implemented to determine prefetching timing is called a prediction confidence model (PCM).
Existing PCMs are modeled to use confidence scores for predicting the subsequent user utterance at the character or word level. However, when considering the actual system response generation, it is important to capture the intent of the user utterance and prepare an appropriate response in advance rather than perfectly predicting every word of the user utterance [5]. In other words, prefetching success depends on whether the user's intent is captured and an appropriate response is prepared.
@@
In constructing our PCM, we investigated the relationship between the predicted and complete user utterances in terms of semantic similarity, based on BERT scores and the success rate of prefetching. We compared our PCM with existing prefetching systems that operate on the word level matching between predicted and complete user utterances. We examined our framework using task-oriented dialogue (TOD) benchmarks: MultiWOZ [6], Spoken-MultiWOZ (SpokenWOZ) [7], and Japanese MultiWOZ (JMultiWOZ) [8]. These different benchmarks were used because PCMs may show different behaviors depending on language structure and the head position of the language. We experimented using gold transcriptions rather than the actual ASR hypothesis with errors, assuming that the speech recognition accuracy was sufficiently high.
@@ 2.1 Problem Definition
Personalized Predictive ASR [9] proposed a pipeline method that prefetches a spoken response and reduces user-perceived latency (UPL). The definition of UPL is shown in Equation (1).
@@ 3 PCM Based on Semantic Similarity
where S-BERT is the Sentence BERT (stsb-xlm-r-multilingual) [12] and $T$ is a threshold.
Our PCM was constructed by following labels defined in the experiment to fine-tune the CLS vector of BERT (bert-base-multilingual-uncased) [13]. We confirmed the following two points in the experiment for this redefinition.
@@ 4.1 Dataset
m
The English and Japanese prediction models were Qwen (Qwen2.5-14B-Instruct) [14] trained on MultiWOZ and JMultiWOZ, respectively. These have been fine-tuned using Low-Rank Adaptation (LoRA) [15] on the user utterances of validation datasets. The inputs are the belief state, up to 4 past dialogue utterances, partial utterances, and a response example. Utterances that do not have dialogue history are not used because they become difficult to predict. At this time, the hyperparameters are as follows: epochs are 1, the learning rate is $2e\text{-}4$, the batch size is 32, the LoRA rank is 16, the optimization function is Adam 8 bit, the maximum number of input tokens is 2048, and the temperature is 1.
@@ 4.3 Training
The PCMs for English and Japanese were fine-tuned on the user utterances (50 dialogues) from the test datasets of MultiWOZ, SpokenWOZ, and JMultiWOZ, respectively. The input features for the PCMs are $h_{\mathrm{dialogue}}$, $\hat{y}_{t}$, and $\hat{y}_{\mathrm{full},t}$. The hyperparameters used for training the PCM are as follows: epoch is $1$, the learning rate is 2e52e-5, the batch size is $16$, the loss function is Focal Loss ($\gamma=2.0$) [16], and the output layer is the Softmax function.
@@ 4.4.1 Automatic Evaluation
The machine evaluation of the prefetched system response group was performed using the following five indicators. Athena-RR [17] was used as the response ranking model.
@@
Since Athena-RR only supports English, the input for evaluating the Japanese PCM was translated into English using GPT3.5 Turbo [18].
@@ 4.4.2 Human Evaluation
We confirmed that prefetched responses were not out of context in dialogue, an essential requirement for dialogue. We conducted the human evaluation of the English and Japanese responses in the prefetched system response group using the following procedure. For Japanese response evaluation, three native Japanese subjects were recruited. For English response evaluation, nine subjects with English proficiency of CEFR B2 [19] or higher were recruited. Each Japanese and English sample was evaluated by three subjects. The experimental procedure is outlined below.

### v2

@@ 1 Introduction
With the progress of research into speech recognition and dialogue response generation, cascade-type spoken dialogue systems (SDSs) are becoming more widely used. Cascade-based SDSs generate spoken responses following two steps [ji2024wavchat]. An automatic speech recognition (ASR) model transcribes a user's speech. A dialogue model generates a response from the transcription. In this sequence, the SDSs generate the response after end-of-sentence/speech (EOS), which is determined by the voice activity detection (VAD) model of ASR. However, this framework causes response delays because the user must wait for the system to generate a response after the user's utterance has ended, which decreases the naturalness of an SDS [jacoby2024human].
Speech response prefetching, which prepares dialogue responses before a user has finished speaking, has been proposed to solve this problem [Kamioka2024response, Zink2024predictive]. If prefetching is successful, the system can respond immediately after detecting EOS. One of the current response prefetching procedures assumes incremental speech recognition and uses a language model to predict complete user utterances from partially recognized utterances. The language model used to predict complete user utterances is called the prediction model. It addresses the delay problem by preparing the system's speech response in advance for the predicted complete user utterance. The system uses a language model to predict the confidence score for the predicted complete utterance to determine the prefetching timing. The language model implemented to determine prefetching timing is called a prediction confidence model (PCM).
Existing PCMs are modeled to use confidence scores for predicting the subsequent user utterance at the character or word level. However, when considering the actual system response generation, it is important to capture the intent of the user utterance and prepare an appropriate response in advance rather than perfectly predicting every word of the user utterance [skantze2005exploring]. In other words, prefetching success depends on whether the user's intent is captured and an appropriate response is prepared.
@@
In constructing our PCM, we investigated the relationship between the predicted and complete user utterances in terms of semantic similarity, based on BERT scores and the success rate of prefetching. We compared our PCM with existing prefetching systems that operate on the word level matching between predicted and complete user utterances. We examined our framework using task-oriented dialogue (TOD) benchmarks: MultiWOZ [zang2020multiwoz], Spoken-MultiWOZ (SpokenWOZ) [si2024spokenwoz], and Japanese MultiWOZ (JMultiWOZ) [ohashi2024jmultiwoz]. These different benchmarks were used because PCMs may show different behaviors depending on language structure and the head position of the language. We experimented using gold transcriptions rather than the actual ASR hypothesis with errors, assuming that the speech recognition accuracy was sufficiently high.
@@ 2.1 Problem Definition
Personalized Predictive ASR [schwarz2023personalized] proposed a pipeline method that prefetches a spoken response and reduces user-perceived latency (UPL). The definition of UPL is shown in Equation (1).
@@ 3 PCM Based on Semantic Similarity
where S-BERT is the Sentence BERT (stsb-xlm-r-multilingual) [reimers-2019-sentence-bert] and $T$ is a threshold.
Our PCM was constructed by following labels defined in the experiment to fine-tune the CLS vector of BERT (bert-base-multilingual-uncased) [DBLP]. We confirmed the following two points in the experiment for this redefinition.
@@ 4.1 Dataset
The English and Japanese prediction models were Qwen (Qwen2.5-14B-Instruct) [qwen2.5] trained on MultiWOZ and JMultiWOZ, respectively. These have been fine-tuned using Low-Rank Adaptation (LoRA) [hu2021lora] on the user utterances of validation datasets. The inputs are the belief state, up to 4 past dialogue utterances, partial utterances, and a response example. Utterances that do not have dialogue history are not used because they become difficult to predict. At this time, the hyperparameters are as follows: epochs are 1, the learning rate is $2e\text{-}4$, the batch size is 32, the LoRA rank is 16, the optimization function is Adam 8 bit, the maximum number of input tokens is 2048, and the temperature is 1.
@@ 4.3 Training
The PCMs for English and Japanese were fine-tuned on the user utterances (50 dialogues) from the test datasets of MultiWOZ, SpokenWOZ, and JMultiWOZ, respectively. The input features for the PCMs are $h_{\mathrm{dialogue}}$, $\hat{y}_{t}$, and $\hat{y}_{\mathrm{full},t}$. The hyperparameters used for training the PCM are as follows: epoch is $1$, the learning rate is 2e52e-5, the batch size is $16$, the loss function is Focal Loss ($\gamma=2.0$) [ross2017focal], and the output layer is the Softmax function.
@@ 4.4.1 Automatic Evaluation
The machine evaluation of the prefetched system response group was performed using the following five indicators. Athena-RR [harrison2023transformer] was used as the response ranking model.
@@
Since Athena-RR only supports English, the input for evaluating the Japanese PCM was translated into English using GPT3.5 Turbo [brown2020language].
@@ 4.4.2 Human Evaluation
We confirmed that prefetched responses were not out of context in dialogue, an essential requirement for dialogue. We conducted the human evaluation of the English and Japanese responses in the prefetched system response group using the following procedure. For Japanese response evaluation, three native Japanese subjects were recruited. For English response evaluation, five subjects with English proficiency of CEFR B2 [council2001common] or higher were recruited. Each Japanese and English sample was evaluated by three subjects. The experimental procedure is outlined below.